OpenStudy (anonymous):

If the distance covered by an object in time t is given by s(t) = t2 + 5t , where s(t) is in meters and t is in seconds, what is the distance covered in the interval between 1 second and 5 seconds? 24 meters 30 meters 40 meters 42 meters 44 meters

OpenStudy (sdfgsdfgs):

s(5) - s(1)

OpenStudy (anonymous):

?

OpenStudy (sdfgsdfgs):

s(1) = distance traveled at 1 second s(5) = distance traveled at 5 seconds So distance traveled between 1 and 5 seconds = s(5) - s(1)

OpenStudy (anonymous):

4 seconds

OpenStudy (sdfgsdfgs):

4 seconds is the time difference; not distance. s(t) = t2 + 5t

OpenStudy (anonymous):

so how do I find the disstance

OpenStudy (sdfgsdfgs):

Calculate s(1) and s(5) then s(5) - s(1)

OpenStudy (anonymous):

44 meters?

OpenStudy (anonymous):

I got a 6 and then I got a 50 and then I subtracted and got a 44

OpenStudy (sdfgsdfgs):

Correct :)

OpenStudy (anonymous):

ok

OpenStudy (wolf1728):

Just to make sure, the equation you are using is d = t*t + 5t